OpenStudy (anonymous):

Input the equation of the given line in slope-intercept form. The line through (1, 0) and parallel to x - y = 7.

OpenStudy (anonymous):

Since its parallel, use the same slope as the given equation and then plug in the point for x and y to solve for b

OpenStudy (anonymous):

-y = -x + 7 y = x - 7 Slope = 1 So, the parallel line also has slope 1. y = x-1